\begin{thebibliography}{20}
\bibitem{AW:AI} A. R. K. Winter, 
 \emph{The Singularity Is Near: When Humans Transcend Biology}, 
 Viking, New York, 2005. 

\bibitem{AW:goldrush} A. C. C. Winter, \emph{The Gold Rush}, 
 United Artists, Hollywood CA, 1925 \&{} 1942. 

\bibitem{AW:tulip} A. C. M. Winter, 
 \emph{Extraordinary Popular Delusions and the Madness of Crowds},
 Richard Bentley, London, 1841. 

\bibitem{AW-1} A. Winter,
 \emph{Mesmerized: Powers of Mind in Victorian Britain},
 The University of Chicago Press, Chicago London, 1998. 

\bibitem{AW0} A. Winter, 
 ``Atypical polymyalgia rheumatica in a 70-year-old male chiropractic patient: a case report'',
 \emph{Clinical Chiropractic}, vol. 7, no. 4, pp. 174-179, 2004.

\bibitem{AW1} A. M. Winter, 
 \emph{Connecting Rural Students to School Using Culturally Relevant Teaching}, 
 doctoral thesis, Johns Hopkins University, 2023. 

\bibitem{AW2} A. Winter, 
 ``Age stratigraphy in the East Antarctic Ice Sheet inferred from radio-echo sounding horizons'', 
  \emph{Earth System Science Data}, vol. 11, no. 3, pp. 1069-1081, 2019. 

\bibitem{AW3} A. Winter, ``Seele als Problem in der Transzendentalphilosophie 
 Kants unter Besonderer Ber\"ucksichtigung des Paralogismus-Kapitels'', 
 in: \emph{Seele: Ihre Wirklichkeit, ihr Verh\"altnis zum Leib und zur menschlichen Person}, 
 K. Kremer (ed.), 
 Series: Studien zur Problemgeschichte der antiken und mittelalterlichen Philosophie, 
 vol. 10, pp. 100-168, E. J. Brill, Leiden K\"oln, 1984. 

\bibitem{AW4} A. Winter \emph{et al.},
 ``Self-assembly of $\pi$-conjugated bis (terpyridine) ligands with zinc (II) ions: 
  new metallosupramolecular materials for optoelectronic applications'', 
  \emph{J. Polymer Science A: Polymer Chemistry}, 
  vol. 47, no. 16, pp. 4083-4098, 2009. 

\bibitem{AW5} A. Winter, 
 ``A genuinely natural information measure'', 
 e-print arXiv [cs.IT]:2103.16662, 2021. 

\bibitem{AW6} A. Winter, 
 \emph{Systems Identification, Flight Performance, and Control of a Fixed-Wing UAV in Icing Conditions}, 
 MSc thesis, Norwegian University of Science and Technology, Department of Engineering
 Cybernetics, and University of Stuttgart, Institute for Flight Mechanics and Control, 2019. 

\bibitem{AW7} A. Winter, 
 ``The Future of Medical Informatics'', 
 \emph{Methods of Information in Medicine}, vol. 48, no. 01, pp. 62-65, 2009. 

\bibitem{AW7.1} A. Winter \emph{et al.}, 
 ``Microwave Hyperthermia for Brain Tumors'',
 \emph{Neurosurgery}, vol. 17, no. 3, pp. 387-399, 1985.

\bibitem{AW7.2} A. Winter, 
 ``Quantum and Classical Message Identification via Quantum Channels'', 
 in: \emph{Festschrift `Alexander S. Holevo 60'} (O. Hirota, ed.), 
 pp. 171-188, Rinton Press Inc., New Jersey, 2004. 

\bibitem{AW8} A. F. de Winter \emph{et al.}, 
 ``Longitudinal patterns and predictors of multiple health risk behaviors 
 among adolescents: The TRAILS study'',
 \emph{Preventive Medicine}, vol. 84, pp. 76-82, 2016.

\bibitem{AW9} A. C. Winters, 
 ``Subseasonal prediction of the state and evolution of the North Pacific jet stream'',
 \emph{Journal of Geophysical Research: Atmospheres}, 
 vol. 126, e2021JD035094, 2021.

\bibitem{AW10} A. Winterbottom \emph{et  al.}, 
 ``Does narrative information bias individual's decision making? A systematic review'', 
 \emph{Social Science \&{} Medicine}, 
 vol. 67, no. 12, pp. 2079-2088, 2008. 

\bibitem{AW11} A. Winter, 
 \emph{Analysis of Uncertainty Estimation Methods to Detect Drifts in Data Streams},
 BSc thesis, TU Darmstadt, 2023. 

\bibitem{AW12} A. R. Winter, 
 ``\,`Faith in the Process': Reading Elly Danica's Don't: A Woman's Word'',
 \emph{Essays on Canadian Writing}, vol. 60, pp. 187-198, Winter 1996.

\bibitem{AW13} J. A. Winter, 
 ``Are Your Troubles Psychosomatic?'', 
 \emph{The Journal of Nervous and Mental Disease}, vol. 120, no. 1, p. 145, 1954.

\bibitem{AW14} A. Winter, 
 ``Die Beurtheilung der Qualit\"at der Frauenmilch nach ihrem mikroskopischen Bilde'',
 \emph{Deutsche Medizinische Wochenschrift}, vol. 28, no. 26, pp. 470-471, 1902. 

\bibitem{AW15} A. D. Winter, 
 ``Distribution Technology and Industrial Planning'', 
 \emph{Royal Australian Planning Institute Journal}, 
 vol. 4, no. 4, pp. 205-209, 1966. 

\bibitem{AW16} A. H. Winter, 
 ``Making a bad calculation'', 
 \emph{Nature Chemistry}, vol. 7, pp. 473-475, 2015. 

\bibitem{AW17} A. T. Winter, 
 ``A model for the fatigue of copper at low plastic strain amplitudes'',
 \emph{The Philosophical Magazine: A Journal of Theoretical Experimental and Applied Physics}, 
 vol. 30, no. 4, pp. 719-738, 1974.

\bibitem{AW18} A. B. Winter, 
 ``Nebraska Home Rule: The Record and Some Recommendations'',
 \emph{Nebraska Law Review}, vol. 59, no. 3, pp. 601-630, 1980.

\bibitem{AW19} J. A. Winter Etcheberry, 
 ``Situaciones actuales en la frontera del principio de culpabilidad'',
 \emph{Revista de Estudios de la Justicia}, vol. 17, pp. 105-143, 2012. 

\bibitem{AW20} A. von Winter, 
 \emph{Beschreibung einer Schlagader-Verletzung in der rechten Ellenbuge Ihrer Majest\"at der 
 K\"oniginn von Baiern, als Folge einer Aderl\"asse und derselben Heilung durch Druckverband,
 nebst einigen darauf Bezug habenden Erl\"auterungen und Reflexionen},
 J. E. von Seidel Kunst- und Buchhandlung, Sulzbach, 1825.

\bibitem{AW21} A. Winters, ``Space Station fluid resupply'',
 in: \emph{NASA, Washington, Technology for Space Station Evolution}, 
 vol. 3 (EVA/Manned Systems/Fluid Management System), pp. 225-240, 1990. 

\bibitem{AW22} A. A. Winter, ``Women's Clubs To-day'', 
 \emph{The North American Review}, vol. 214, no. 792, pp. 636-640, 1921. 

\bibitem{AW23} A. Winter, 
 ``On environment-assisted capacities of quantum channels'', 
 \emph{Markov Processes Related Fields}, vol. 13, no. 1-2, pp. 297-314, 2007.

\bibitem{AW24} A. Winter \emph{et al.}, 
 ``Quantum Phase Transition in the Sub-Ohmic Spin-Boson Model: 
 Quantum Monte Carlo Study with a Continuous Imaginary Time Cluster Algorithm'',
 \emph{Physical Review Letters}, vol. 102, no. 3, 030601, 2009. 

\bibitem{AW25} N. A. Winter, 
 ``Il tetto di Caprifico a confronto'', 
 \emph{Archeologia Classica}, vol. 57, pp. 530-535, 2006.

\bibitem{AW26} A. Winter \emph{et al.}, 
 ``Poleward expansion of the coccolithophore \textit{Emiliania huxleyi}'',
 \emph{J. Plankton Research}, vol. 36, no. 2, pp. 316-325, 2014.

\bibitem{AW27} R. A. Winter, 
 ``Vertical control and price versus nonprice competition'',
 \emph{The Quarterly Journal of Economics}, vol. 108, no. 1, pp. 61-75, 1993.

\bibitem{AW28} A. M. Winter, 
 ``Social services: effective practices in serving survivors of torture'',
 \emph{Quarterly Journal on Rehabilitation of Torture Victims and Prevention of Torture},
 vol. 21, no. 1, pp. 48-55, 2011. 

\bibitem{AW29} A. McLean Winter, 
 ``The Laughing Doves of 1812 and the Satiric Endowment of Antiwar 
 Rhetoric in the United States'',
 \emph{Publications of the Modern Language Association of America}, 
 vol. 124, no. 5, pp. 1562-1581, 2009. 

\bibitem{AW:justiz} A. F. D. Winter, 
 \emph{Justiz}, Diogenes Verlag, Z\"urich, 1985. 

\bibitem{AW30} A. Winter Noble, 
 ``Em Tempos de Reformas, a Alfab\'etica'', 
 \emph{Paralelo 31}, vol. 2, no. 13, pp. 214-232, 2019.

\bibitem{AW31} A. Winterer, 
 \emph{Verkehrte Sexualit\"at -- ein umstrittenes Pauluswort}, 
 doctoral dissertation, European University Studies, vol. 810, 
 Frankfurt am Main, 2005. 

\bibitem{AW32} A. Winterest, 
 ``Sempre pi\`u bello'', \emph{Bella Figura}, vol. 1, no. 1, pp. 10-16, 2019.

\end{thebibliography}

\documentclass[conference,a4paper]{IEEEtran}
\usepackage{bbding}
\usepackage[T1]{fontenc}
\usepackage{aurical}
\usepackage{huncial}
\usepackage{linearb}
\usepackage{textcomp}
\usepackage{amssymb,epic,eepic}
\usepackage{amsmath}
\usepackage[amsmath,amsthm,thmmarks]{ntheorem}
\usepackage{amsfonts}
\usepackage{stmaryrd}
\usepackage{capt-of}
\usepackage{graphicx}
\usepackage[usenames,dvipsnames]{pstricks}
\usepackage{epsfig}
\usepackage{pst-grad}
\usepackage{pst-plot}
\usepackage{mathtools}
\usepackage{textfit} 
\usepackage{subfigure}
\usepackage{float}

\usepackage{comment}

\newtheorem{theorem}{Theorem} 

\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{conjecture}[theorem]{Conjecture}

\begin{document}

\title{Is Winter Coming?}

\author{\IEEEauthorblockN{A. Winter}
\IEEEauthorblockA{\textit{Grup d'Informaci\'o Qu\`antica}\\
\textit{Departament de F\'isica}\\
\textit{Universitat Aut\`onoma}\\
\textit{de Barcelona, Spain}\\
andreas.winter@uab.cat} 

\and
\IEEEauthorblockN{A. Winter}
\IEEEauthorblockA{\textit{ICREA--Instituci\'{o}}\\
\textit{Catalana de Recerca}\\
\textit{i Estudis Avan\c{c}ats}\\
\textit{Barcelona, Spain}}

\and
\IEEEauthorblockN{A. Winter}
\IEEEauthorblockA{\textit{Institute for}\\
\textit{Advanced Study}\\
\textit{Technische Universit\"at}\\
\textit{M\"unchen, Germany}} 

\and
\IEEEauthorblockN{A. Winter}
\IEEEauthorblockA{\textit{QUIRCK--Quantum}\\
\textit{Independent Research}\\ 
\textit{Center Kessenich}\\
\textit{Bonn, Germany}}
}

\maketitle


\begin{abstract}
We critically examine the often-made observation that 
``quantum winter [or some other winter] is coming'', and the related 
admonition to prepare for this or that winter, inevitably bound to arrive. 
What we find based on even the most superficial look at the available 
evidence is that such statements not only are overblown hype, but are also 
factually wrong: Winter is here, and the real question is rather for how 
long it/they will stay.
\end{abstract}

\begin{IEEEkeywords}
Winter, wintry, winterise, wintering.
\end{IEEEkeywords}

\medskip\noindent
\textit{\textbf{1. Introduction.}}
It has been suggested, and is now being repeated often without proper
attribution, that ``quantum winter is coming''. Based on the familiar 
seasonal cycle, such a statement might be considered so banal as to be 
almost empty, even if referring metaphorically to the so-called 
``hype cycle'' of fashionable and overly-promoted fields of research. 
Quantum winter in this sense is a source of considerable anxiety 
among members of the currently objectively overhyped field of quantum 
science and technology. The field of artificial intelligence has gone 
through several cycles of hype and disillusionment, and indeed AI winter, 
the original ``winter'' of what is now recognised as a collection of similar 
phenomena across different fields of scientific research and/or technological 
innovation, has been proclaimed repeatedly since the 1980s \cite{AW:AI}. 

Rather obviously, a lot of this talk is merely the product of minds 
intoxicated by the rush of attention (both socially and monetary), 
and at the same time dreading heavily the end of the halcyon days of easy 
living off the hype. 
Quite characteristically, there is usually a secondary investment winter
affecting those financing the hype and rightly fearing for the expected 
return on their investment. 
The whole thing is not unlike the gold rushes of another age, when 
desperate, crazed diggers and ruthless entrepreneurs turned entire 
regions upside down, only to move on when the fever had passed and 
the ground did not yield any more of the precious metal \cite{AW:goldrush}. 
The market pressure exerted by modern-day investors sometimes, paradoxically, 
prolongs the hype phase, preventing the drop into the wintry 
valley of disillusionment, as derivatives are created (spin-off 
companies floated on the stock market, etc) whose speculation value 
can carry the bubble for some time when its practitioners are already 
preparing for the crash. 
It is precisely this -- partial or repressed -- knowledge of a bubble ripe 
for bursting that creates the broadly shared dread of the ``winter'' to come. 
Note that economically driven hypes are nothing recent in Western economies, 
as evidenced by the famous tulip mania in the Dutch 17th century \cite{AW:tulip}
(cf.~\cite{AW-1}). 

While we are personally sympathetic to the psychological conditions affecting 
our dear colleagues, as scientists it is our duty to regard the matter with 
the cold eye of empiricism. As we will explain in the next two 
sections, the fear of any winter coming is entirely unfounded, 
principally because winter, in all its forms and manifestations, 
has long arrived. What conclusions can be drawn from this fact 
is discussed in the final section.

\medskip\noindent
\textit{\textbf{2. Methods.}}
We follow the model of \cite{AW0}, elaborated in \cite{AW1,AW2,AW3,AW4} and subsequent 
investigations, to estimate the proximity of winter. Our method is 
scientific and self-referential, i.e.~bibliographic. 
We need a small auxiliary result. 

\begin{lemma}[cf.~A.~Winter~\cite{AW5}]
\label{lemma:Winter}
There exists an integer $n$, significantly smaller than one might believe, 
such that the reader of a scientific article, when confronted with $n$ 
or more arbitrary references, will accept anything the paper 
claims to follow from them.
\end{lemma}
\begin{proof}\textit{(Sketch)}
To be added in the journal version, see however \cite{AW6,AW7} and 
\cite{AW7.1,AW7.2} for similar statements. 
\end{proof}

\medskip\noindent
\textit{\textbf{3. Winter's universality.}}
With these surprisingly elementary preparations we are ready for 
our main result. It is enough to regard 
Refs.~\cite{AW8,AW9,AW10,AW11,AW12,AW13,AW14,AW15}, 
in particular \cite{AW16}, followed by the litany of 
\cite{AW17,AW18,AW19,AW20,AW21,AW22,AW23,AW24,AW25,AW26,AW27,AW28,AW29}, 
and to consider that this is only a small and sufficiently random sample 
of wintry \emph{selecta}, to realise beyond doubt that Winter has been 
around quite forcefully for decades, if not centuries. 

This clearly means that any fear of Winter coming is completely 
irrational: there is nothing to be prepared about, to ``winterise'' 
as it were, as we are right in the middle of it. 
We allow ourselves the expression of the modest hope that this 
insight will lead to the acceptance of the universal presence of Winter 
in all walks of (academic) life \cite{AW:justiz}.

\medskip\noindent
\textit{\textbf{4. Epilogue.}}
We regard our arguments as irrefutable. Hence, as a next step, we need 
to prepare for the phase after acceptance: the hype. That there are 
already indications of a wintering crescendo is evident from the 
literature \cite{AW30,AW31,AW32}.
So, will a Winter winter come, and should we worry about it? 

More to the point, shall we compare something to a summer's day? 
Of course, one swallow does not a summer make. 
However, inspired by these expressions of somewhat complementary sentiments, 
and taking the theme to its natural conclusion, 
we are motivated to make the following conjecture. 

\begin{conjecture}
\label{conj:winter}
Now is the Winter of our discontent made glorious summer by this sun of York. 
\end{conjecture}

This might appear at first sight rather bold, and will perhaps be 
suspected of being indebted more to wishful thinking than having 
any basis in solid evidence. We invite the interested reader to 
try and falsify it, as we have tried -- and failed. In the appendix, 
A. Winter graciously shares her extensive remarks on possible approaches 
towards this conjecture, and on interesting consequences of it.

\medskip\noindent
\textit{\textbf{Acknowledgments.}}
The authors thank A. Winter and A. Winter for their thoughtful and 
interesting comments, in particular with regard to A. Winter's contribution
to the field. 
They are grateful to A. Winter for spotting an error in an early version 
of Lemma \ref{lemma:Winter}. Finally, they acknowledge A. Winter's 
thorough proofreading. 
The first (AW) and fourth (AW) author were supported by the 
Andrew Andreevich and Andrea Winter Foundation through a generous 
research stipend. 
A. Winter declares no competing financial interests.

 
\medskip\noindent
\textit{\textbf{Appendix --- Remarks on Conjecture \ref{conj:winter} (by A. Winter).}}
This was to be something profound reflecting my complex and nuanced
opinions about the provocative work of A. Winter \emph{et al.}, 
but the present paragraph is too short to contain it. 
For

\end{document}